princeevee:

Given that sinθ=3/5 , what is cosθ?

Hero:

So in this case, the same rules apply except all you need to know in this case is that \(\sin(\theta) = \dfrac{a}{c}\) and \(\cos(\theta) = \dfrac{b}{c}\) and Pythagorean Theorem: \(a^2 + b^2 = c^2\)

princeevee:

@Hero

princeevee:

@Vocaloid

Hero:

That doesn't apply here @princeevee

Hero:

You only need to use what I have given you above. The other stuff only applied to other problem. Doesn't apply here except for what is above.

princeevee:

so it's given that a=3 and b=5?

Hero:

actually a = 3 and c = 5

princeevee:

oh right, sorry, so a is 3, and b is x as of now

Hero:

or rather b is "unknown"

princeevee:

c^2 = 25 since 5^2, and a is 9

princeevee:

25 - 9 is 16

princeevee:

so b^2 is 16, i think

princeevee:

but it isnt actually B as of yet, right?

princeevee:

@Hero

Hero:

Nope, because there's still something left to do.

princeevee:

what's that?

princeevee:

subtraction?

Hero:

if \(b^2 = 16\) what do you need to do in order to isolate b?

princeevee:

well i just counted in my head for that, and got 4

princeevee:

meaning that cos(theta) is 4/5?

Hero:

Yes correct, but you don't have to count in your head. That takes forever

princeevee:

sorry

princeevee:

think you could help with more?